Sarah is twice as old as her youngest brother. If the difference between their ages is 15 years, how old is her youngest brother?

Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

Sarah is twice as old as her youngest brother. If the difference between their ages is 15 years, how old is her youngest brother?
---------
Equations:
s = 2b
s-b = 15 yrs
-----
Substitute for "s" and solve for "b"::
2b - b = 15
b = 15 yrs (brother's age now)
-------------
Cheers,
Stan H.
-----------

Answer by addingup(3677)

If Sarah is 2times as old as her brother, then 2*brother = Sarah
S= 2b We are going to use this value for Sarah
S-b= 15 In this formula, plug in the value above for Sarah:
2b-b= 15
b= 15 And Sarah, the problem says, is twice as old:
15*2= 30 Sarah is 30. And the problem says Sarah is 15 yrs older:
30-15= 15 We have the correct answer.